OpenStudy (anonymous):

You are having a conference call with the CEO of a paper company. You have interpreted the number of trees cut down versus profit as the function P(x) = -x4 + x3 + 7x2 - x - 6. Describe to the CEO what the graph looks like. Use complete sentences, and focus on the end behaviors of the graph and where the company will break even (where P(x) = 0). help pleaseeee!!!

OpenStudy (anonymous):

@mathmale

OpenStudy (anonymous):

@SolomonZelman

OpenStudy (anonymous):

@undeadknight26

OpenStudy (solomonzelman):

Okay, so the graph of P(x) would be, this https://www.desmos.com/calculator/dsk5edlkpy And seeming, when x→ ±∞ then y→ -∞ (whether x → ∞ or x→ - ∞ )

OpenStudy (solomonzelman):

p(x)=0, you will need to set \(\normalsize\color{blue}{ P(x)=-x^4+x^3+7x^2-x-6}\) \(\normalsize\color{blue}{ 0=-x^4+x^3+7x^2-x-6}\) and sovle for x.

OpenStudy (anonymous):

and how would you solve for x in the problem?

OpenStudy (solomonzelman):

I will give you a hint: \(\normalsize\color{blue}{ -x^4+x^3+7x^2-x-6=-(x-1)(x+1)(x+2)(x+3)}\)

OpenStudy (solomonzelman):

So, \(\normalsize\color{blue}{-x^4+x^3+7x^2-x-6=0}\) \(\normalsize\color{blue}{-(x-1)(x+1)(x+2)(x+3)=0}\) \(\normalsize\color{blue}{(x-1)(x+1)(x+2)(x+3)=0}\) the roots are ?

OpenStudy (solomonzelman):

roots, x-intercepts, same thing ;)

OpenStudy (anonymous):

arent they 1,2,3,6 ?

OpenStudy (solomonzelman):

\(\normalsize\color{blue}{(x-1)(x+1)(x+2)(x+3)=0}\) \(\huge\color{blue}{~↓~↓~~↓~↓ }\) 1 -1 -2 -3

OpenStudy (solomonzelman):

Makes sense ?

OpenStudy (anonymous):

okay so the roots are 1,-1 ,-2,-3?

OpenStudy (solomonzelman):

yes

OpenStudy (anonymous):

so what do i do next

OpenStudy (anonymous):

are you still there ? @SolomonZelman